What is the final velocity of the proton in the previous problem? (The mass of the proton is 1.67 × 10–27 kg.) A. 8 × 1014 m/s B. 1,000 m/s C. 3 × 108 m/s D. 3 × 105 m/s
Answer #1

A)

Kinetic energy= q*v

=e*500V

=500 eV

Ans: B

B)

Kinetic energy=500 eV

=500*1.6*10^-19 J

= 8*10^-17 J

Kinetic energy is also given as 0.5 *m*v^2

so,

0.5 *m*v^2 = 8*10^-17

0.5*(1.67*10^-27)*v^2=8*10^-17 J

v^2=9*10^10

v=3*10^5 m/s

Ans: D